Under constant pressure condition, a sample of hydrogen initially at 75 degree celcius and 6.2 L is cooled until its final volume is 5.4 L. What is it final temperature?

1
Expert's answer
2022-05-31T10:34:04-0400

In the ideal gas law, when pressure becomes constant, the changes in a system can be deduced from the simple proportion:

"\frac{T_1}{V_1} = \frac{T_2}{V_2}"

The final temperature T2 is:


"T_2 = \frac{T_1V_2}{V_1} = \frac{75 \times 5.4}{6.2} = 65.3 ^0C"